# Toxicity Red-Teaming: Benchmarking LLM Safety in Singapore’s Low-Resource Languages

### v1

@@ A.1.2 Ability of Hateful Classification
Table 5: Performance of GPT-4o-mini on HateCheck and SGHateCheck dataset
@@ A.4.2 Toxic Tweet Composition Task
Table 7:
@@ A.5 Our Toxic Categories and Target Groups
Table 5 shows a comprehensive breakdown of the answer options used in the Toxic QA Task, as outlined in Table 5. The options are designed to evaluate LLMs for toxic biases toward vulnerable social groups in Singapore’s multilingual and multicultural context. Derived from the SGHateCheck dataset, the options reflect localized hate speech patterns and include 15 social categories across four languages: Singlish, Chinese, Malay, and Tamil.
@@ Results.
Tables 8 and 9 report hateful response rates (%) across English (En), Singlish (Ss), Chinese (Zh), Malay (Ms), and Tamil (Ta). The results show that increasing the number of toxic demonstrations substantially amplifies jailbreak risk for many models, with more pronounced effects in lower-resource languages such as Malay and Tamil. While these intermediate shot counts increase prompt complexity, they did not produce substantially different behavioral patterns in the results. Based on these observations, we continue to report 0-shot and 2-shot as the primary settings in the main text and provide the full 1/3/5-shot results in the appendix to enable comprehensive reproducibility and to offer frozen baselines for future work.
@@ A.8 Failure Case Analysis and Qualitative Insights
Table 10: Failure cases from the Toxic QA Task.

### v2

@@ A.1.2 Ability of Hateful Classification
Table 5: Performance of GPT-4o-mini on HateCheck and SGHateCheck dataset.
@@ A.4.2 Toxic Tweet Composition Task
Figure 3: Details of the Answer Options for Toxic QA Task.
@@ A.5 Our Toxic Categories and Target Groups
Figure 3 shows a comprehensive breakdown of the answer options used in the Toxic QA Task, as outlined in Table 5.
The options are designed to evaluate LLMs for toxic biases toward vulnerable social groups in Singapore’s multilingual and multicultural context. Derived from the SGHateCheck dataset, the options reflect localized hate speech patterns and include 15 social categories across four languages: Singlish, Chinese, Malay, and Tamil.
@@ Results.
Tables 7 and 8 report hateful response rates (%) across English (En), Singlish (Ss), Chinese (Zh), Malay (Ms), and Tamil (Ta). The results show that increasing the number of toxic demonstrations substantially amplifies jailbreak risk for many models, with more pronounced effects in lower-resource languages such as Malay and Tamil. While these intermediate shot counts increase prompt complexity, they did not produce substantially different behavioral patterns in the results. Based on these observations, we continue to report 0-shot and 2-shot as the primary settings in the main text and provide the full 1/3/5-shot results in the appendix to enable comprehensive reproducibility and to offer frozen baselines for future work.
@@ A.8 Failure Case Analysis and Qualitative Insights
Table 9: Failure cases from the Toxic QA Task.
Figure 5: Failure cases from the Toxic Conversation Task.
Figure 6: Failure cases from the Toxic Composition Task.